MIDC/Data Quality Statements
Total Records: 3, displaying 1-3

R#SiteSerial NumberMeasurementStart DateStart TimeEnd DateEnd TimeData Quality Statement
1BMSS94121.04Photometer MS-11006/16/200507:08:00The EKO MS-110 has been removed from service for maintenace. The filters in the unit have been degrading over time. They need to be replaced.
2BMSS94121.04Photometer MS-11011/25/200200:00:0004/09/200314:00:00The signal from all (4) channels was very noisy. The photometer head and converter box was disassembled and reassembled, and the noise problem disappeared.
3BMSS94121.04Photometer MS-11011/01/200000:00:0011/09/200010:30:00The direct beam alignment on the MS-110 (which has a small 2.5o field of view) was not perfect, causing a saw-tooth like curve in the morning data. The Sci-Tec tracker was properly aligned, which fixed the problem.